I've tried 0.16, 0.48, 0.36

and 32, 96, 72

I have 2 trials left, please help.

Here's a hint: Hint 2. How to calculate the expected frequencies of a different example population Consider an example population of individuals that have two alleles for a specific locus, AB and AC. In the population, 70% (0.7) of the alleles are AB, and 30% (0.3) of the alleles are AC. The expected frequencies of each genotype can be calculated using the equation for Hardy-Weinberg equilibrium (p^2+2pq+q^2=1). The expected frequency of the ABAB genotype is p2=0.72=0.49. The expected frequency of the ABAC genotype is 2pq=2(0.7)(0.3)=0.42. The expected frequency of the ACAC genotype is q2=0.32=0.09. Notice that the expected frequencies of the three genotypes add up to 1. p^2+2pq+q^2=1 0.49+0.42+0.09=1

Part B Determining the expected frequency of each genotype Considering the same population of cats as in Part A, what is the expected frequency of each genotype (T-TSTSTS) based on the equation for Hardy-Weinberg equilibrium? Keep in mind that you just learned in Part A that The allele frequency of T is 0.4 The allele frequency of T is 0.6 The equation for Hardy-Weinberg equilibrium states that at a locus with two alleles, as in this cat population, the three genotypes will occur in specific proportions: Enter the values for the expected frequency of each genotype:TT Weinberg equation to this cat population, see Hints 1 and 2. TTS, and TSTS Enter your answers numerically to two decimal places, not as percentages. For help applying the Hardy Hints reset? help Phenotype Genotype(tail length) Number of individuals in population Expected frequency TTL long 60 036 0.16 TLTS medium 40 0.48 TSTS short 100 Submit My Answers Give Up Incorrect; Try Again; 3 attempts remaining
Answer #1

Answer:

TLTL = (q^2= q * q) 0.4 * 0.4 = 0.16

TLTS =(2pq) 2 * 0.4 * 0.6 =0.48

TSTS = (p^2 = p*p) = 0.6 * 0.6 = 0.36